Use method of undetermined coefficients method to solve the following

y''+y=5xsinx,y(0)=0,y'(0)=1


1
Expert's answer
2022-03-24T15:21:35-0400

y''+y=5xsinx,y(0)=0,y'(0)=1

The solutions to a nonhomogeneous equation are of the form

"y(x) = y_c(x) + y_p(x)" ,

where "y_c" is the general solution to the associated homogeneous equation and "y_p" is a particular solution.

The associated homogeneous equation:

"y''+y=0"

The general solution of this equation is determined by the roots of the characteristic equation:

"\lambda^2+1=0"

"\lambda=\pm i" , where "i=\sqrt{-1}" .

"y_c(x)=c_1\cos{x}+c_2\sin{x}" .

The particular solution of the differential equation:

"y_p(x)=x((Ax+B)\sin{x}+(Cx+D)\cos{x})"

"y_p'=(2 A x + B - C x^2 - D x)\sin{x} +""(A x^2 + B x + 2 C x + D)\cos{x}"

"y_p''= (-A x^2 + 2 A - B x - 4 C x - 2 D)\sin{x} +""(4 A x + 2 B - C x^2 + 2 C - D x)\cos{x}"

Now put these into the original differential equation to get:

"(-A x^2 + 2 A - B x - 4 C x - 2 D)\sin{x} +""(4 A x + 2 B - C x^2 + 2 C - D x)\cos{x} +""x((Ax+B)\sin{x}+(Cx+D)\cos{x})=""5x\sin{x}"

Equating coefficients, we get

"A=0"; "D=0"; "C=-\frac54"; "B=\frac54".

"y_p(x)=\frac54x(\sin{x}-x\cos{x})"

"y=c_1\cos{x}+c_2\sin{x}+\frac54x(\sin{x}-x\cos{x})"

"y(0)=c_1=0"

"y\u2019(0)=c_2=1"

Answer:

"y=\sin{x}+\frac54x(\sin{x}-x\cos{x})" .